Fermi National Laboratory
 

Accelerator Update for November 9, 2007 - November 12, 2007

 

Friday November 9

The day shift began with the Tevatron (TeV), Booster, Linac and many other areas off due to an empty nitrogen DEWAR at TeV sector A0 used to keep vacuum and LCW valves open, with Operations monitoring the Recycler stash (191.8E10) and the Antiproton Source (Pbar) stack (172.2E10), with the Main Injector (MI) waiting on the other accelerators, with the Collider Detector at Fermilab (CDF), D-Zero (D0), the Meson MTest, and SciBooNE / MiniBooNE experiments waiting for beam.

Operations reported at 8:25 AM that Pbar LCW pressures were recovering. The Cryo System Coordinator diverted one of their nitrogen deliveries to go to TeV sector A0 and fill the DEWAR.

Operations prepared Linac and Booster at 8:49 AM for an access to turn on vacuum pumps and other small but important jobs.

An expert reported at 9:21 AM that CUB had stabilized due to a return of nitrogen pressure.

Operations began preparing TeV F-sector and A-E sectors for access at 12:46 PM.

Operations had CDF and D0 prepared for access by 1 PM.

CDF came out of access at 1:55 PM.

At 3:22 PM, a Linac expert reported that a Linac RF station Power Amplifier (LRF5) had failed and needed to be replaced.

Operations reported at 5:21 PM that they were putting the TeV through the first of three HEP ramps.

At 5:49 PM, the 750keV area buncher tripped off. It reset.

Linac experts reported at 5:53 PM that they had completed their repairs on LRF5.

At 6:03 PM, the TeV suffered a quench at sectors A4, B1, BC, and BD.

Operations established beam to Linac, Booster, and Sci/MiniBooNE at 6:28 PM.

At 6:52 PM, LRF5 tripped off on a high cavity pressure indication. It reset.

At 9:21 PM, TRF4 tripped off. An operator contacted a RF expert and then reset the station.

At 9:53 PM, the MiniBooNE horn tripped off. It reset.

At 9:54 PM, The TeV suffered a quench at BA and BB.

The MTest experiment stopped using beam at midnight.

Saturday November 10

The midnight shift began with Operations monitoring the stash (188.9E10) and the stack (165.2E10), with the TeV recovering from a quench, and with Sci /MiniBooNE taking beam.

At 12:48 AM, Operations reported that vacuum pump power supply for the H- Source had tripped off. It reset.

Operations put the TeV through a dry squeeze at 1:15 AM.

At 1:27 AM, Operations reported that a TeV quadrupole power supply (C:D0Q5) had tripped off during the dry squeeze. It reset.

At 2:11 AM, LRF5 tripped off. It reset.

At 6:03 AM, a Linac buncher tripped off. It reset.

Operations established beam to the MTest experiment at 7:40 AM.

Operations put the TeV into shot setup at 10:15 AM.

At 12:38 PM, Operations reported that a TeV power supply (T:VE12) had tripped off on an over temperature indication. Operators replaced its regulator chassis. This and a problem with an antiproton emittance blowup had put a halt to the TeV shot setup.

Operations put the TeV back into shot setup at 1:05 PM.

At 2:27 PM, LRF5 tripped off. It reset.

Operations established store 5705 at 4:09 PM with an initial luminosity of 33.0E30.

At 4:25 PM, a separator at TeV sector C49 sparked and caused a 25% loss of luminosity.

At 6:49 PM, the TeV suffered a quench and sector F4. The CDF SVX detector pulled the abort due to bad beam positions. Operators investigated and reported that the initial problem was probably due to a sector B11V separator spark. Their investigations continued.

Operations reported at 7:23 PM that both CDF and D0 said that their new collision points were good.

Operations reported at 8 PM that the Recycler had stashed antiprotons once during the last 12 hours.

At 8:58 PM, Operations reported that a TeV power supply (S:TF19) had tripped off for the first of four times in about 11 hours. It reset each time.

TeV experts put the TeV through the first of a number of dry squeezes at 9:13 PM to work on antiproton beam positions during the ramp.

At 11:48 PM, the Linac low energy buncher tripped off. It reset.

MTest reported that they were through with the beam for the night at midnight.

Sunday November 11

The midnight shift began with TeV experts working on the ramp, Operations monitoring the stash, with Pbar stacking, and with Sci /MiniBooNE taking beam.

At 2:19 AM, the Electron Cooling tripped off and the Pelletron would not reset itself. Operations contacted an expert and, with his help, managed to get everything reset. Later, another expert called in and had operators make an adjustment that would allow the Pelletron to do its own normal reset.

Operations put the TeV into shot setup at 2:32 AM.

At 3:59 AM, a Booster critical device controller tripped off and would not reset. Operators called in interlock experts.

At 4:16 AM, Operations reported having problems with TeV collimators.

At 4:33 AM, the TeV suffered a quench at sectors A4, B1, and F4. Later, operators reported that the probable cause was due to the sector F48 collimator.

Operations reported that Booster beam had been restored by 5:47 AM. Interlock experts reported finding and replacing a blown fuse.

Operations established beam to MTest at 7:45 AM.

Operations reported at 8 AM that the Recycler had stashed once during the last 12 hours.

A TeV expert arrived at 9:38 AM to further diagnose the TeV quench. He confirmed that it was a collimator problem.

At 9:59 AM, S:VTF19 tripped off and would not reset. Operators contacted an expert and then changed the regulator chassis. It still would not reset. An expert came in and changed the control chassis.

Operations put the TeV into shot setup at 10:56 AM.

Operations, with the help of TeV and Recycler experts, established store 5714 at 1:13 PM with an initial luminosity of 62.4E30.

At 3:39 PM, Operations reported a slight luminosity drop at CDF and D0. Everything looked normal and no cause was found.

At 6:10 PM, the TeV luminosity took another small dip. Operators found that T:RFST3 had tripped off. Operators turned off the station and its water heaters.

Operations reported at 6:58 PM that a Booster expert began a corrector magnet study.

Operations reported at 8 PM that the Recycler had stashed once during the last 12 hours.

At 9:17 PM, the 750 KeV area buncher tripped off. It reset.

A TeV expert made some final adjustments at 10:07 PM to the D0 collision point move.

Monday November 12

The midnight shift began with Operations monitoring store 5714 and the stash, with Pbar stacking, and with MTest and Sci /MiniBooNE taking beam.

MTest reported at 12:09 AM that they were through with the beam for the night.

At 2:09 AM, Operations reported that they received a level 1 alarm for the TeV sector A0 nitrogen DEWAR. An expert reported that we didn’t get a scheduled delivery on Sunday. A delivery will supposedly come later this afternoon. The expert will try to get them to come in earlier.

At 5:08 AM, the 750 KeV area buncher tripped off. It reset.

Operations established beam to MTest at 6:10 AM.

Operations terminated store 5714 at 8:02 AM.

The Plans for Monday

The plans for today are to complete the TeV cryo lead work, to stack, stash, and to establish a store. The Run Coordinator will allow the accesses required by the new antiproton source experiment, T972. They will take data during the day and then make plans for tomorrow.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 11/12/2007 -->    email Fermilab